OpenStudy (anonymous):

A force of 73.8 newtons is applied to travel bag at an angle 39.0° with the horizontal. What are the horizontal and vertical components of the force?

OpenStudy (anonymous):

 using\[\cos(\theta)=F_x/F\] and \[\sin(\theta)=F_y/F\] where F is the resultant force (73.8 N) you find the components by: \[F_x=73.8*\cos(39.0)\]\[F_y=73.8*\sin(39.0)\]